# A Survey on Diffusion Models for Anomaly Detection

### v2

@@ 1 Introduction
The ever-increasing volume, velocity, and variety of data present both opportunities and challenges for anomaly detection (AD).  In cybersecurity, real-time threat detection is crucial given the constant influx of network traffic and logs Liu et al. (2024b).  Similarly, robust fraud detection systems are essential for financial institutions processing massive transaction datasets Wang et al. (2023a).  Healthcare and manufacturing also rely heavily on data for early disease diagnosis and predictive maintenance, respectively Zhang et al. (2023). Such applications highlight the growing need for automated AD methods to efficiently identify outliers in complex datasets. However, traditional techniques, often based on statistical methods or rule-based systems, struggle with the scale and complexity of modern data Katsuoka et al. (2024).  Specifically, these methods often require extensive manual feature engineering and struggle to adapt to evolving data distributions.  Additionally, the sheer data volume can overwhelm traditional methods, rendering them computationally infeasible for real-time applications. Consequently, more sophisticated and scalable AD approaches are needed. For example, weakly supervised approaches leverage limited labeled data to improve performance Liu et al. (2024c), driving innovation toward scalable deep learning architectures that address these fundamental challenges.
@@
Recent growth in complex and multi-dimensional data has created new challenges for anomaly detection methods. In this context, DMs have emerged as a promising solution due to their inherent connection to density estimation Le Lan and Dinh (2021). DMs learn the probability distributions of normal data through iterative denoising. Leveraging their exceptional generative abilities, DMs accurately reconstruct normal patterns while capturing underlying manifold structures, enabling anomaly detection through reconstruction error analysis and probability density estimation Hu and Jin (2023).  For example, AnoDDPM Wyatt et al. (2022) uses diffusion models for anomaly detection via reconstruction error, and ODD Wang et al. (2023b) enhances this by leveraging advanced generative modeling for high-dimensional data.  Additionally, the learned score function, representing the gradient of the log-probability density, can be directly used as an anomaly score, as demonstrated in diffusion time estimation Livernoche et al. (2023).
@@
While existing surveys have separately reviewed anomaly detection Liu et al. (2024b, c) and diffusion models Luo (2023); Yang et al. (2024b), they either focus on traditional deep learning approaches without considering DMs’ emerging role, or treat anomaly detection merely as one of many DM applications without in-depth analysis. The field of diffusion models for anomaly detection (DMAD) still lacks a systematic and comprehensive review that specifically examines how DMs enhance anomaly detection capabilities, which motivates our work to provide a comprehensive and timely overview of DMAD, capturing its current state and future developments.
@@ 2.1 Anomaly Detection
Anomaly detection aims to identify instances deviating significantly from the expected data distribution Wang and Vastola (2023). Given a dataset $\mathcal{D}=\{x_{1},x_{2},...,x_{n}\}$, where $x_{i}\in\mathbb{R}^{d}$, the goal is to identify the anomalous subset $\mathcal{A}\subset\mathcal{D}$.  Anomalous behavior is characterized by low probability density under the typically unknown distribution $p(x)$.  Several anomaly types exist, including point anomalies Hu and Jin (2023) that are individual deviant instances; contextual anomalies Sui et al. (2024) that are anomalous within a specific context; and collective anomalies that are groups of instances anomalous only when considered together. Common evaluation metrics include AUC Xu et al. (2023), measuring the classifier’s ability to distinguish between normal and anomalous instances.  Additionally, precision, recall, and F1-score assess the trade-off between correct identification and minimizing false alarms Wang et al. (2023a); Xiao et al. (2023); Zuo et al. (2024). The choice of metric depends on the application and the relative importance of minimizing false positives and negatives Liu et al. (2024c).
@@ 2.2 Diffusion Models
Forward process is known as the diffusion process, gradually adds noise to data $x_{0}$ over $T$ timesteps, transforming its distribution into an isotropic Gaussian. Specifically, given a variance schedule $\{\beta_{t}\}_{t=1}^{T}$, where $10<\beta_{1}<\dots<\beta_{T}<$ 1, the process generates sequence of latent variables $\{x_{1},\dots,x_{T}\}$ via a Markov chain with the Gaussian transition kernel:
@@
Here, $\mathcal{N}(x;\mu,\Sigma)$ denotes a Gaussian distribution with mean $\mu$ and covariance $\Sigma$, and $\mathbf{I}$ is the identity matrix. The parameter $\beta_{t}$ controls the noise added at each step.  As $t$ increases, $x_{T}$ approaches Gaussian noise. Due to the Markov property, the joint distribution is:
@@
where $\mu_{\theta}$ and $\Sigma_{\theta}$ are the predicted mean and covariance, respectively, parameterized by a neural network (e.g., U-Net Yan et al. (2024) or Transformer Cao et al. (2024)) and conditioned on the noisy input $x_{t}$ at timestep $t$.  While the covariance $\Sigma_{\theta}(x_{t},t)$ can be learned, a common simplification is to use a fixed, time-dependent variance Salimans and Ho (2021). The neural network learns to predict the mean $\mu_{\theta}(x_{t},t)$, which guides denoising. Several parameterization strategies exist, including predicting the mean of $q(x_{t-1}|x_{t},x_{0})$, added noise $\epsilon_{t}$, or original data $x_{0}$. Starting from Gaussian noise $x_{T}$, the reverse process iteratively applies $p_{\theta}(x_{t-1}|x_{t})$ until $x_{0}$ is reached.
@@
DMs are trained by maximizing the likelihood of observed data.  However, due to the latent variable formulation, direct optimization is intractable. Consequently, training relies on maximizing a variational lower bound (VLB) of the data log-likelihood, derived using variational inference Bercea et al. (2023) as:
@@
where $q$ and $p_{\theta}$ denote the forward and learned reverse process distributions, respectively.  Here, $D_{KL}$ is the Kullback-Leibler divergence, $x_{0}$ represents the original data, and $x_{t}$ is the noisy sample at timestep $t$. A simplified objective minimizes the $L_{2}$ distance between true and predicted noise Han et al. (2024b), can be expressed as:
@@ 2.3 Score-based Models and Score Matching
Recent research highlights the inherent linear structure within score-based models, particularly at higher noise levels, potentially accelerating sampling Wang and Vastola (2023).  Prioritizing specific noise levels during training, based on perceptual relevance, can further improve performance Choi et al. (2022).  In addition, gradient guidance, incorporating an external objective function’s gradient during sampling, adapts pre-trained models to specific tasks Tur et al. (2023b).
@@ 2.4 Variants of Diffusion Models
DDPMs Bercea et al. (2023) progressively introduce Gaussian noise to the data across multiple time steps using a Markov chain. A trained neural network then reverses this process, enabling the model to effectively learn complex distributions, which makes it particularly well-suited for anomaly detection Wyatt et al. (2022). However, generating samples can be computationally expensive.
@@
Score SDEs Deveney et al. (2023) model diffusion as a continuous-time SDE, where the score function, representing the gradient of the log-probability density, can be used for anomaly scoring Li et al. (2024). While offering flexibility for complex distributions Livernoche et al. (2023), training and parameterization choices can be challenging Graham et al. (2023).
@@ 3.1 Reconstruction-based Anomaly Detection
Figure 2: Illustration of reconstruction-based AD methods: (a) basic, (b) latent space, and (c) conditional reconstructions.
@@
RAD hinges on the principle that DMs trained on normal data reconstruct normal instances effectively, but struggle with anomalous ones Hu and Jin (2023).  Specifically, after training on normal samples, the model reconstructs a test sample by reversing the diffusion process. The difference between the input and its reconstruction constitutes the anomaly score Zhang et al. (2023). A higher score indicates a greater likelihood of an anomaly.  For instance, GLAD Yao et al. (2025) employs global and local reconstruction perspectives.  In contrast, AutoDDPM Bercea et al. (2023) focuses on improving normal region reconstruction.  Alternatively, some approaches use a learned similarity network for comparison, yielding a semantic distance as the anomaly score Wang et al. (2023b).
@@
RAD leverages latent space representations for enhanced efficiency and effective handling of high-dimensional data, such as images or videos. Specifically, projecting input data into a lower-dimensional latent space significantly reduces the computational burden associated with reconstruction. Autoencoders (AEs) are commonly employed for this dimensionality reduction, which learns a compressed representation by encoding data into a latent space and then decoding it back to the original space.  Integrating AEs with DMs offers a promising approach. For example, DMs trained on AE-learned latent representations can capture the normal data distribution in the compressed space Hu and Jin (2023).  In addition, NGLS-Diff Han et al. (2024a) is a novel approach that utilizes DMs within a normal gathering latent space to enhance anomaly detection capabilities for time series data.  The latent space reconstruction error then serves as an anomaly score, with deviations indicating potential anomalies Wang et al. (2023c). Shared latent spaces across multiple AEs can separate treatment and context.  Similarly, supervised dimensionality reduction can optimize the latent space for specific classification tasks, and employing Brownian motion within Diffusion VAEs captures dataset topology Le Lan and Dinh (2021).
@@ 3.2 Density-based Anomaly Detection
Figure 3: Illustration of density-based AD methods: (a) score function-based and (b) diffusion time estimation methods.
@@ 4.1 Image Anomaly Detection
DISYRE Naval Marimont et al. (2024) C CamCAN, ATLAS, BraTS AP: 0.45 (ATLAS), 0.51 (BraTS-T1), 0.73 (BraTS-T2), etc.
@@
Table 1: Summary of IAD methods for metric results on datasets.
IAD represents a key application area for DMs, focusing on identifying deviations from normality in both global patterns affecting entire images and local anomalies confined to specific regions. Many approaches leverage DMs’ reconstruction capabilities, with methods calculating anomaly scores from reconstruction errors after rebuilding images from noisy inputs Xu et al. (2023); Tur et al. (2023a); Wang et al. (2023d). Advanced frameworks like GLAD Yao et al. (2025) employ global and local adaptive approaches, while ODD Wang et al. (2023b) introduces similarity networks for semantic distance measurement. Medical imaging applications have shown particular promise, with FNDM Li et al. (2023) and mDDPM Iqbal et al. (2024) demonstrating enhanced efficiency through masked diffusion models, while Fontanella et al. (2024) introduces novel counterfactual generation techniques. Recent research includes AnomalyDiffusion Hu et al. (2024b) for few-shot generation and Naval Marimont et al. (2024)’s cold-diffusion approach. MDPS Wu et al. (2024) and DRAD Sheng et al. (2024) further enhance detection through masked sampling and noise embedding, while methods like Tebbe and Tayyub (2024) employ dynamic step size computation to improve localization accuracy.
@@ 4.2 Time Series Anomaly Detection
SimpDM Liu et al. (2024d) C Iris, Yacht, etc. RMSE: 0.059 (Iris), 0.045 (Yacht), 0.102 (Housing), etc.
CoDi Lee et al. (2023a) C & D Bank, Heart, etc. B-F1: 0.4726, B-AUROC: 0.8106, M-F1: 0.62211, etc., on Bank
TimeDiT Cao et al. (2024) C Stocks, Energy, ect. DS: 0.115 (Stocks), 0.1778 (Air Quality), 0.1726 (Energy), etc.
@@
Table 2: Summary of TSAD methods for metric results on datasets.
DMs have emerged as a powerful technique for TSAD, effectively learning complex distributions of sequential data Yang et al. (2024b). DMs can identify various anomaly types, including point, contextual, and collective anomalies Sui et al. (2024). Advanced approaches like NGLS-Diff Han et al. (2024a) leverage latent spaces, while methods Zuo et al. (2024) used structured state space layers explicitly to model temporal dependencies. Recently, some innovations address key challenges: D3R Wang et al. (2023a) tackles non-stationarity through dynamic decomposition, while TimeDiT Cao et al. (2024) introduces a foundation model approach with multiple masking schemes. Another key challenge is non-stationarity. DiffAD Xiao et al. (2023) handle missing data through imputation, complemented by self-supervised techniques Liu et al. (2024d) and co-evolving strategies Lee et al. (2023a) for mixed-type temporal data.
@@ 4.3 Video Anomaly Detection
Table 3: Summary of VAD methods for metric results on datasets.
@@ 4.4 Multimodal Anomaly Detection
MPDR Yoon et al. (2023) I, V, A C MNIST, CIFAR-10, etc. AUPR: 0.764 (MNIST), 0.9860 (CIFAR-10), 0.8338 (CIFAR-100), etc.
@@
AnomalyXFusion Hu et al. (2024a) I, T, Te C MVTec-AD, MVTec LOCO IS: 1.82, IC-LPIPS: 0.33
Notes: Mods. indicates modalities, where ’I’ is Image, ’V’ is Vector, ’A’ is Audio, ’T’ is Text, and ’Te’ is Texture)
Table 4: Summary of MAD methods for metric results on datasets.

### v5

@@ 1 Introduction
The ever-increasing volume, velocity, and variety of data present both opportunities and challenges for anomaly detection (AD).  In cybersecurity, real-time threat detection is crucial given the constant influx of network traffic and logs Jin et al. (2024).  Similarly, robust fraud detection systems are essential for financial institutions processing massive transaction datasets Wang et al. (2023a).  Healthcare and manufacturing also rely heavily on data for early disease diagnosis and predictive maintenance, respectively Zhang et al. (2023). Such applications highlight the growing need for automated AD methods to efficiently identify outliers in complex datasets. However, traditional techniques, often based on statistical methods or rule-based systems, struggle with the scale and complexity of modern data Katsuoka et al. (2024).  Specifically, these methods often require extensive manual feature engineering and struggle to adapt to evolving data distributions.  Additionally, the sheer data volume can overwhelm traditional methods, rendering them computationally infeasible for real-time applications. Consequently, more sophisticated and scalable AD approaches are needed. For example, weakly supervised approaches leverage limited labeled data to improve performance, driving innovation toward scalable deep learning architectures that address these fundamental challenges Pang et al. (2021).
@@
Recent growth in complex and multi-dimensional data has created new challenges for anomaly detection methods. In this context, DMs have emerged as a promising solution due to their inherent connection to density estimation Le Lan and Dinh (2021). DMs learn the probability distributions of normal data through iterative denoising. Leveraging their exceptional generative abilities, DMs accurately reconstruct normal patterns while capturing underlying manifold structures, enabling anomaly detection through reconstruction error analysis and probability density estimation Hu and Jin (2023); Salimans and Ho (2021).  For example, AnoDDPM Wyatt et al. (2022) uses diffusion models for anomaly detection via reconstruction error, and ODD Wang et al. (2023b) enhances this by leveraging advanced generative modeling for high-dimensional data.  Additionally, the learned score function, representing the gradient of the log-probability density, can be directly used as an anomaly score, as demonstrated in diffusion time estimation Livernoche et al. (2023).
@@
While existing surveys have separately reviewed anomaly detection Jin et al. (2024); Pang et al. (2021) and diffusion models Yang et al. (2024b); Luo (2023), they either focus on traditional deep learning approaches without considering the emerging role of diffusion models, or treat anomaly detection merely as one of many DM applications without in-depth analysis. Recent breakthroughs in diffusion models for anomaly detection (DMAD) demonstrate remarkable potential to revolutionize the field through their unique capabilities in modeling complex data distributions and generating high-quality samples. Despite these advances, there is no existing work has systematically reviewed how DMs enhance anomaly detection or analyzed their emerging role at the intersection of both domains. Our work addresses this gap by providing a timely overview of DMAD, analyzing its theoretical foundations, current achievements, and promising future directions that could fundamentally reshape the landscape of anomaly detection research.
@@ 2.1 Anomaly Detection
Anomaly detection aims to identify instances deviating significantly from the expected data distribution Wang and Vastola (2023). Given a dataset $\mathcal{D}=\{x_{1},x_{2},...,x_{n}\}$, where $x_{i}\in\mathbb{R}^{d}$, the goal is to identify the anomalous subset $\mathcal{A}\subset\mathcal{D}$.  Anomalous behavior is characterized by low probability density under the typically unknown distribution $p(x)$.  Several anomaly types exist, including point anomalies Hu and Jin (2023) that are individual deviant instances; contextual anomalies Sui et al. (2024) that are anomalous within a specific context; and collective anomalies that are groups of instances anomalous only when considered together. Common evaluation metrics include AUC Xu et al. (2023), measuring the classifier’s ability to distinguish between normal and anomalous instances.  Additionally, precision, recall, and F1-score assess the trade-off between correct identification and minimizing false alarms Wang et al. (2023a); Xiao et al. (2023); Zuo et al. (2024). The choice of metric depends on the application and the relative importance of minimizing false positives and negatives Pang et al. (2021).
@@ 2.2 Diffusion Models
Forward process is known as the diffusion process, gradually adds noise to data $x_{0}$ over $T$ timesteps, transforming its distribution into an isotropic Gaussian Wyatt et al. (2022). Specifically, given a variance schedule $\{\beta_{t}\}_{t=1}^{T}$, where $10<\beta_{1}<\dots<\beta_{T}<$ 1, the process generates sequence of latent variables $\{x_{1},\dots,x_{T}\}$ via a Markov chain with the Gaussian transition kernel:
@@
Here, $\mathcal{N}(x;\mu,\Sigma)$ denotes a Gaussian distribution with mean $\mu$ and covariance $\Sigma$, and $\mathbf{I}$ is the identity matrix. The parameter $\beta_{t}$ controls the noise added at each step.  As $t$ increases, $x_{T}$ approaches Gaussian noise. Due to the Markov property Li et al. (2023), the joint distribution is:
@@
where $\mu_{\theta}$ and $\Sigma_{\theta}$ are the predicted mean and covariance, respectively, parameterized by a carefully designed neural network (e.g., U-Net Yan et al. (2024) or Transformer Cao et al. (2024)) and conditioned on the noisy input $x_{t}$ at timestep $t$.  While the covariance $\Sigma_{\theta}(x_{t},t)$ can be learned, a common simplification is to use a fixed, time-dependent variance Salimans and Ho (2021). The neural network learns to effectively predict the mean $\mu_{\theta}(x_{t},t)$, which guides denoising. Several parameterization strategies exist, including predicting the mean of $q(x_{t-1}|x_{t},x_{0})$, added noise $\epsilon_{t}$, or original data $x_{0}$. Starting from Gaussian noise $x_{T}$, the reverse process iteratively applies $p_{\theta}(x_{t-1}|x_{t})$ until $x_{0}$ is reached.
@@
DMs are trained by maximizing the likelihood of observed data.  However, due to the complex latent variable formulation, direct optimization is computationally intractable. Consequently, training relies on maximizing a variational lower bound (VLB) of the data log-likelihood, derived using variational inference Bercea et al. (2023) as:
@@
where $q$ and $p_{\theta}$ denote the forward and learned reverse process distributions, respectively.  Here, $D_{KL}$ is the standard Kullback-Leibler divergence, $x_{0}$ represents the original data, and $x_{t}$ is the noisy sample at timestep $t$. A simplified objective function minimizes the $L_{2}$ distance between true and predicted noise Han et al. (2024b), can be expressed as:
@@ 2.3 Score-based Models and Score Matching
Recent research highlights the inherent linear structure within score-based models, particularly at higher noise levels, potentially accelerating sampling Wang and Vastola (2023).  Prioritizing selected specific noise levels during training, based on perceptual relevance, can further improve performance Choi et al. (2022).  In addition, gradient guidance, incorporating an external objective function’s gradient during sampling, effectively adapts pre-trained models to specific tasks Tur et al. (2023b).
@@ 2.4 Variants of Diffusion Models
DDPMs Bercea et al. (2023) progressively introduce controlled Gaussian noise to the data across multiple time steps using a Markov chain. A trained neural network then reverses this process, enabling the model to effectively learn complex distributions, which makes it particularly well-suited for anomaly detection Wyatt et al. (2022). However, generating samples can be computationally expensive.
@@
Score SDEs Deveney et al. (2023) model diffusion as a continuous-time SDE, where the score function, representing the gradient of the log-probability density, can be effectively used for anomaly scoring Li et al. (2024). While offering flexibility for complex distributions Livernoche et al. (2023), training and parameterization choices can be particularly challenging Graham et al. (2023).
@@ 3 Methodologies
Recent research efforts on DMAD can be broadly categorized into three primary methodologies: reconstruction-based, density-based, and hybrid approaches.
@@ 3.1 Reconstruction-based Anomaly Detection
Figure 2: Pipeline illustration of reconstruction-based AD methods: (a) basic, (b) latent space, and (c) conditional reconstruction.
@@
RAD hinges on the fundamental principle that DMs trained on normal data reconstruct normal instances effectively, but struggle with anomalous ones Hu and Jin (2023).  Specifically, after training on normal samples, the model reconstructs a test sample by reversing the diffusion process. The difference between the input and its reconstruction constitutes the anomaly score Zhang et al. (2023). A higher score indicates a greater likelihood of an anomaly.  For instance, GLAD Yao et al. (2025) employs global and local reconstruction perspectives.  In contrast, AutoDDPM Bercea et al. (2023) focuses on improving normal region reconstruction.  Alternatively, some approaches use a learned similarity network for comparison, yielding a semantic distance as the anomaly score Wang et al. (2023b).
@@
RAD leverages latent space representations for enhanced efficiency and effective handling of high-dimensional data, such as images or videos. Specifically, projecting input data into a lower-dimensional latent space significantly reduces the computational burden associated with reconstruction. Autoencoders (AEs) are commonly employed for this dimensionality reduction, which learns a compressed representation by encoding data into a latent space and then decoding it back to the original space.  Integrating AEs with DMs offers a promising approach. For example, DMs trained on AEs-learned latent representations can capture the normal data distribution in the compressed space Hu and Jin (2023).  In addition, NGLS-Diff Han et al. (2024a) is a novel approach that utilizes DMs within a normal gathering latent space to enhance anomaly detection capabilities for time series data.  The latent space reconstruction error then serves as an anomaly score, with deviations indicating potential anomalies Wang et al. (2023c). Shared latent spaces across multiple AEs can separate treatment and context.  Similarly, supervised dimensionality reduction can optimize the latent space for specific classification tasks, and employing Brownian motion within Diffusion VAEs captures dataset topology Le Lan and Dinh (2021).
@@ 3.2 Density-based Anomaly Detection
Figure 3: Pipeline illustration of density-based AD methods: (a) score function-based and (b) diffusion time estimation method.
@@ 4.1 Image Anomaly Detection
DISYRE Naval Marimont et al. (2024) C CamCAN, ATLAS, BraTS AP: 45% (ATLAS), 51% (BraTS-T1), 73% (BraTS-T2), etc.
@@
Table 1: Summary of IAD methods with metric results on datasets.
IAD represents a key application area for DMs, focusing on identifying deviations from normality in both global patterns affecting entire images and local anomalies confined to specific regions. Many approaches leverage DMs’ reconstruction capabilities, with methods calculating anomaly scores from reconstruction errors after reconstructing images from noisy inputs Xu et al. (2023); Tur et al. (2023a); Wang et al. (2023d). Advanced frameworks like GLAD Yao et al. (2025) employ global and local adaptive approaches, while ODD Wang et al. (2023b) introduces similarity networks for semantic distance measurement. Medical imaging applications have shown particular promise, with FNDM Li et al. (2023) and mDDPM Iqbal et al. (2024) demonstrating enhanced efficiency through masked diffusion models, while Fontanella et al. (2024) introduces novel counterfactual generation techniques. Recent research includes AnomalyDiffusion Hu et al. (2024b) for few-shot generation and DISYRE Naval Marimont et al. (2024) for synthetic anomaly generation in medical imaging.. MDPS Wu et al. (2024) and DRAD Sheng et al. (2024) further enhance detection through masked sampling and noise embedding, while methods like DIC Tebbe and Tayyub (2024) employ dynamic step size computation to improve localization accuracy.
@@ 4.2 Time Series Anomaly Detection
SimpDM Liu et al. (2024b) C Iris, Yacht, etc. RMSE: 0.059 (Iris), 0.045 (Yacht), 0.102 (Housing), etc.
CoDi Lee et al. (2023a) C & D Bank, Heart, etc. B-F1: 47.26%, B-AUROC: 81.06%, M-F1: 62.21%, etc., on Bank
TimeDiT Cao et al. (2024) C Stocks, Energy, ect. DS: 11.50% (Stocks), 17.78% (Air Quality), 17.26% (Energy), etc.
@@
Table 2: Summary of TSAD methods with metric results on datasets.
DMs have emerged as a powerful technique for TSAD, effectively learning complex distributions of sequential data Yang et al. (2024b). DMs can identify various anomaly types, including point, contextual, and collective anomalies Sui et al. (2024). Advanced approaches like NGLS-Diff Han et al. (2024a) leverage latent spaces, while methods Zuo et al. (2024) used structured state space layers explicitly to model temporal dependencies. Recently, some innovations address key challenges: D3R Wang et al. (2023a) tackles non-stationarity through dynamic decomposition, while TimeDiT Cao et al. (2024) introduces a foundation model approach with multiple masking schemes. Another key challenge is non-stationarity. DiffAD Xiao et al. (2023) handle missing data through imputation, complemented by self-supervised techniques Liu et al. (2024b) and co-evolving strategies Lee et al. (2023a) for mixed-type temporal data.
@@ 4.3 Video Anomaly Detection
Table 3: Summary of VAD methods with metric results on datasets.
@@ 4.4 Multimodal Anomaly Detection
MPDR Yoon et al. (2023) I, V, A C MNIST, CIFAR-10, etc. AUPR: 76.40% (MNIST), 98.60% (CIFAR-10), 83.38% (CIFAR-100), etc.
@@
AnomalyXFusion Hu et al. (2024a) I, T, Te C MVTec-AD, MVTec LOCO IS: 1.82 (MVTec-AD), IC-LPIPS: 0.33 (MVTec-AD), etc
Notes: Mods. indicates modalities, where ’I’ is image, ’V’ is vector, ’A’ is audio, ’T’ is text, and ’Te’ is texture.
Table 4: Summary of MAD methods with metric results on datasets.